OpenStudy (anonymous):

hello, how can use differentials to approximate the possible propagated error in computing the area of the triangle?

OpenStudy (anonymous):

A = (1/2) b h dA = (1/2) h db + (1/2) b dh dA, db, dh are error estimates Make them fractional estimates by using dA/A = etc.